OpenStudy (anonymous):

A ladder leaning against a wall makes a 60o angle with the ground. The base of the ladder is 4 m from the building. How high above the ground is the top of the ladder?

OpenStudy (anonymous):

you have a 60 degree angle so you have a 30-60-90 right triangle 4m is opposite the 30 degree angle so the hyp is 8 [the side opposite 30 degrees is half the hypotenuse ] 8^2=x^2+4^2 64=x^2+16 x^2=48 x=sqrt4=4sqrt3 A regular hexagon is composed of 12 congruent 30-60-90 triangles. If the length of the hypotenuse of one of those triangles is 183, find the perimeter of the hexagon. ? a hexagon has 6 sides you will have to explain how the triangles are set up so the 12 triangles form 6 sides

OpenStudy (anonymous):

Right angled triangle. Base = 4m Angle with ground = 60 degrees Height of top of ladder is found by tan = Opposite/Adjacent tan (60) = Opposite/4 Opposite = 4 + tan(60) = 6.9m This is the height of the top of ladder above ground. Hope this helps. :-)